OpenStudy (lexy148):

how to find inverse function for h(x)=x+5/8

zepdrix (zepdrix):

Hey Lexy! Is your function \(\large\rm h(x)=\frac{x+5}{8}\) or \(\large\rm h(x)=x+\frac{5}{8}\) ? If it's the first one, please try to get in the habit of using brackets! >.< h(x) = (x+5)/8